# Reward-Shifted Speculative Sampling Is An Efficient Test-Time Weak-to-Strong Aligner

### v2

@@ Abstract
Aligning large language models (LLMs) with human preferences has become a critical step in their development. Recent research has increasingly focused on test-time alignment, where additional compute is allocated during inference to enhance LLM safety and reasoning capabilities. However, these test-time alignment techniques often incur substantial inference costs, limiting their practical application. We are inspired by the speculative sampling acceleration, which leverages a small draft model to efficiently predict future tokens, to address the efficiency bottleneck of test-time alignment. We introduce the reward-Shifted Speculative Sampling (SSS) algorithm, in which the draft model is aligned with human preferences, while the target model remains unchanged. We theoretically demonstrate that the distributional shift between the aligned draft model and the unaligned target model can be exploited to recover the RLHF optimal solution without actually obtaining it, by modifying the acceptance criterion and bonus token distribution. Our algorithm achieves superior gold reward scores at a significantly reduced inference cost in test-time weak-to-strong alignment experiments, thereby validating both its effectiveness and efficiency.
@@ 2.2 Speculative Sampling
$\pi_{\text{bonus}}(\cdot|x,y_{<t})=\left(\pi_{\text{ref}}(\cdot|x,y_{<t})-\pi_{\text{draft}}(\cdot|x,y_{<t})\right)_{+},$ (4)
@@ 3.2 Speculative Sampling under Draft-Target Distributional Shift
   if $\epsilon<p_{\text{accept}}(t)$ in Eq. (6then
@@ 4.1 Experimental Settings
For post-training draft models, we use the TRL library [^1] https://github.com/huggingface/trl. for supervised fine-tuning (SFT) and direct preference optimization (DPO) (Rafailov et al., 2023) with the hyper-parameters in Table 3. These hyper-parameter choices are based on the settings of Tao and Li (2025) and adjusted for better performance by grid search.
@@
For inference experiments, we use the Transformers library [^2] https://github.com/huggingface/transformers. with a temperature of 0.8 and a maximum sequence length of 128 tokens. These settings are standard in previous works (Chen et al., 2023; Khanov et al., 2024; Li et al., 2025).
@@ 4.2 Alignment Quality and Inference Efficiency
Table˜4 demonstrates the performance advantages of SSS across different model pairs on HH-RLHF for a single run. We compare against a range of test-time alignment baselines, including Best-of-$N$ (BoN), TreeBoN (Qiu et al., 2024), and CARDS (Li et al., 2025). We keep the reward model sizes used in these baselines the same as draft models for fair comparison. The baseline choices represent a spectrum of test-time alignment strategies balancing alignment quality and computational efficiency. We also include comparisons with a standard speculative sampling (SS) approach.
@@
Table 4: Comparison of test-time weak-to-strong alignment methods in terms of gold reward, number of LLM calls, and inference time. “Gold R” refers to the average score assigned by the gold reward model, “# Calls” indicates the number of forward passes through the target model $\pi_{\text{ref}}$ (a proxy for computational cost), and “Time” is the actual wall-clock inference time per response. “Speedup” is computed relative to the BoN baseline for each method.
@@ A.1 Test-Time Alignment
Early alignment research focused on post-training LLMs with techniques like RLHF (Ouyang et al., 2022), but the high cost of policy optimization methods like PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) has motivated the test-time (or decoding-time) alignment approaches that leave the LLMs’ parameters frozen. For example, best-of-$N$ (BoN) generates multiple complete responses and select the highest-reward one; rejection sampling continues generating proposal responses until a reward threshold is met. Built upon these 2 techniques, test-time alignment methods at varying granularities are developed. For example, token-level reward-guided search (Deng and Raffel, 2023; Khanov et al., 2024), segment-level rejection sampling (Li et al., 2025), and segment-level MC tree search (Qiu et al., 2024) have all been explored. However, although they improve the efficiency of test-time alignment, they still rely on external reward models to interact with the decoding process, which is the major cause of the inefficiency. Our algorithm (SSS) jumps out of the frameworks with reward models, and leverages a small draft model to reflect human preference. The interaction between draft and target models are far more efficient than the traditional LLM-RM framework.
@@ D.3 Is SSS Really Approaching the RLHF Optimal Solution?
   if $\epsilon<p_{\text{accept}}(t)$ in Eq. (3then

### v3

@@ Abstract
Aligning large language models (LLMs) with human preferences has become a critical step in their development. Recent research has increasingly focused on test-time alignment, where additional compute is allocated during inference to enhance LLM safety and reasoning capabilities. However, these test-time alignment techniques often incur substantial inference costs, limiting their practical application. We are inspired by the speculative sampling acceleration, which leverages a small draft model to efficiently predict future tokens, to address the efficiency bottleneck of test-time alignment. We introduce the reward-Shifted Speculative Sampling (SSS) algorithm, in which the draft model is aligned with human preferences, while the target model remains unchanged. We theoretically demonstrate that the distributional shift between the aligned draft model and the unaligned target model can be exploited to recover the RLHF optimal solution without actually obtaining it, by modifying the acceptance criterion and bonus token distribution. Our algorithm achieves superior gold reward scores at a significantly reduced inference cost in test-time weak-to-strong alignment experiments, thereby validating both its effectiveness and efficiency. [^1] The implementation of this method can be found in our codebase: https://github.com/lblaoke/CARDS.
@@ 2.2 Speculative Sampling
$\pi_{\text{bonus}}(\cdot|x,y_{<t})$ (4)
$=\left(\pi_{\text{ref}}(\cdot|x,y_{<t})-\pi_{\text{draft}}(\cdot|x,y_{<t})\right)_{+},$
@@ 3.2 Speculative Sampling under Draft-Target Distributional Shift
   if $\epsilon<p_{\text{accept}}(t)$ in Eq. (6) then
@@ 4.1 Experimental Settings
For post-training draft models, we use the TRL library (von Werra et al., 2020) for supervised fine-tuning (SFT) and direct preference optimization (DPO) (Rafailov et al., 2023) with the hyper-parameters in Table 3. These hyper-parameter choices are based on the settings of Tao and Li (2025) and adjusted for better performance by grid search.
@@
For inference experiments, we use the Transformers library (Wolf et al., 2020) with a temperature of 0.8 and a maximum sequence length of 128 tokens. These settings are standard in previous works (Chen et al., 2023; Khanov et al., 2024; Li et al., 2025).
@@ 4.2 Alignment Quality and Inference Efficiency
Table˜4 demonstrates the performance advantages of SSS across different model pairs on HH-RLHF for a single run. We compare against a range of test-time alignment baselines, including Best‑of‑$N$ (BoN), TreeBoN (Qiu et al., 2024), and CARDS (Li et al., 2025). We keep the reward model sizes used in these baselines the same as draft models for fair comparison. The baseline choices represent a spectrum of test-time alignment strategies balancing alignment quality and computational efficiency. We also include comparisons with a standard speculative sampling (SS) approach.
@@
Table 4: Comparison of test-time weak-to-strong alignment methods in terms of gold reward, number of LLM calls, and inference time. “Gold R” refers to the average score assigned by the gold reward model, “# Calls” indicates the number of forward passes through the target model $\pi_{\text{ref}}$, and “Time” is the actual wall-clock inference time per response. “Speedup” is computed relative to the BoN baseline for each method.
@@ A.1 Test-Time Alignment
Early alignment research focused on post‑training LLMs with techniques like RLHF (Ouyang et al., 2022), but the high cost of policy optimization methods like PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) has motivated the test‑time (or decoding‑time) alignment approaches that leave the LLMs’ parameters frozen. For example, best-of-$N$ (BoN) generates multiple complete responses and select the highest‑reward one; rejection sampling continues generating proposal responses until a reward threshold is met. Built upon these 2 techniques, test-time alignment methods at varying granularities are developed. For example, token‑level reward‑guided search (Deng and Raffel, 2023; Khanov et al., 2024), segment-level rejection sampling (Li et al., 2025), and segment-level MC tree search (Qiu et al., 2024) have all been explored. However, although they improve the efficiency of test-time alignment, they still rely on external reward models to interact with the decoding process, which is the major cause of the inefficiency. Our algorithm (SSS) jumps out of the frameworks with reward models, and leverages a small draft model to reflect human preference. The interaction between draft and target models are far more efficient than the traditional LLM-RM framework.
@@ D.3 Is SSS Really Approaching the RLHF Optimal Solution?
   if $\epsilon<p_{\text{accept}}(t)$ in Eq. (3) then